# One Communication Round is All It Needs for Federated Fine-Tuning Foundation Models

### v1

@@ Appendix C Additional Experimental Results
Standalone Results of Local Models.
To further demonstrate the effectiveness of federated fine-tuning, we performed the standalone experiment to compare the performance of the global model and the local model only trained on local datasets. We did the experiments on the llama-7b model and Wizard dataset and displayed the results in Table 6. The results show that the accuracy of most local models is slightly lower than that of the global model, with some local models outperforming the global model. This is reasonable in the context of the federated fine-tuning task because the models have already been pre-trained. Therefore, even though clients have less training data, the performance of local models does not differ significantly from the global model.
@@
Table 6: Standalone results of 3-epochs federated fine-tuning on Llama-7b with Wizard dataset. The numeric header columns indicate the ARC Challenge accuracy of the local models only fine-tuned on their local dataset for 3 epochs.
One-Shot 0 1 2 3 4 5 6 7 8 9
51.88 50.79 51.02 50.05 50.43 52.33 51.22 51.28 52.30 51.21 51.11
Figure 7: The MT-bench score of global model in 1-5 global rounds.
More Global Round Settings.
@@
We also tested the model performance when we had more and fewer global rounds in a multi-round setting. We evaluated the global model in 1, 2, 3, 4, and 5 global rounds when fine-tuning the Llama-7b model on Wizard dataset. The results are shown in Figure 7. In the first round, the MT-bench score increases from the 2.86 in base model to around 3.80. Then, it slightly increases towards 3.90 in the 3rd round and begins to decrease afterward. A similar phenomenon can be seen in other datasets and models that the model performance will increase in the initial 2-4 rounds and then gradually decline due to overfitting. Thus, we use 3 global rounds in all of the multi-round experiments.
Appendix D Proof of Theorem 1
According to Eq. 7 and Eq. 8, ignoring the learning rates, the difference of the global model can be bounded by:
Considering Assumption 1, we have:
$\begin{split}\varepsilon\leq\sum_{j=k+1}^{Tk}Lm\|(\bm{w}_{i}^{(0,j)}-\bm{w}_{i}^{(t,j-kt)}\|,\end{split}$ (11)
According to Assumption 2, we can deduce:
$\begin{split}\varepsilon\leq\sum_{j=k+1}^{Tk}L\tau m\|\bm{w}^{(0,0)}\|,\end{split}$ (12)
Thus we have:
$\begin{split}\varepsilon\leq L\tau Tkm\|\bm{w}^{(0,0)}\|,\end{split}$ (13)
which is Theorem 1.
Appendix E Limitation
This work has two main limitations. (1) The paper is limited in federated fine-tuning tasks since we lack the computation resources to conduct federated pre-training experiments. (2) Since common stable diffusion models do not vary significantly in parameter size, this work does not observe the performance of different-sized stable diffusion models in one-shot federated fine-tuning. The impact of model parameter size on one-shot federated fine-tuning in text-to-image generation tasks still needs to be explored.

### v2

@@ IV-B Main Results
Figure 6: Standalone results of one-shot federated fine-tuning on Llama-7B with Wizard dataset. Bars 0–9 are the accuracy of local models with the corresponding index; Global is the one-shot aggregated global model.
Standalone Results of Local Models. To further demonstrate the effectiveness of one-shot federated fine-tuning, we performed the standalone experiment to compare the performance of the global model and the local model only trained on local datasets. We did the experiments on the llama-7b model and Wizard dataset and displayed the results in Fig. 6. The results show that the accuracy of local models is slightly lower than that of the global model, with some local models perform similar to the global model. This is reasonable in the context of the federated fine-tuning task because the models have already been pre-trained. Therefore, even though clients have less training data, the performance of local models does not differ significantly from the global model. This further supports that a single aggregation can already capture most of the attainable gain.
Figure 7: The MT-bench score of the global model in 1–5 global rounds.
@@
More Global Round Settings. We also tested the model performance when we had more and fewer global rounds in a multi-round setting. We evaluated the global model in 1, 2, 3, 4, and 5 global rounds when fine-tuning the Llama-7b model on Wizard dataset. The results are shown in Fig. 7. In the first round, the MT-bench score increases from the 2.86 in base model to around 3.80. Then, it slightly increases towards 3.90 in the 3rd round and begins to decrease afterward. A similar phenomenon can be seen in other datasets and models that the model performance will increase in the initial 2-4 rounds and then gradually decline due to overfitting. Thus, we use 3 global rounds in all of the multi-round experiments.
V Discussion
One-Shot Federated Fine-Tuning Saves Communication Cost
In FL, the server needs to send the model parameters to all the selected clients and receive the model updates from the clients in each communication round [ghiasvand2024communication]. Thus, the total number of communicated parameters in multi-round should be $2mTS$, where $S$ is the model size. In one-shot federated fine-tuning, the server and the clients only perform one-round communication, so the number of communicated parameters is only $2mS$. This reduction in communication overhead is significant, especially in the federated fine-tuning of large FMs. For instance, the Llama-13b model has approximately 50 GB parameters, i.e., $S=50$ GB. In our experiments, the three-round federated fine-tuning on Llama-13b needs to communicate about 3000 GB data between the server and the clients, which may be unaffordable in scenarios with tight communication budgets, whereas one-shot reduces this amount to about 1000 GB. Beyond raw transfer size, fewer rounds simplify orchestration and reduce exposure to transient failures during long jobs, since there are fewer synchronization points and fewer opportunities for timeouts. In practice, one-shot can be combined with standard techniques such as quantization, sparsification, and parameter-efficient tuning so that the effective $S$ is smaller, which further improves communication efficiency without changing the algorithmic interface. This makes the approach easier to deploy in bandwidth-constrained environments and lowers the barrier for cross-organization collaboration where network conditions and maintenance windows are variable.
One-Shot Federated Fine-Tuning Supports Asynchronous Global Aggregation
In traditional multi-round FL, clients need to train local models synchronously. The server can only perform the aggregation and send the new global model to clients after receiving all local model updates. This requirement poses challenges for federated learning applications. For example, if local computation resources are occupied by other tasks or if the connection between the server and clients is unstable, the training process will be halted. One-shot federated fine-tuning effectively addresses this problem. The server can update the global model with local updates as soon as they are received, allowing for real-time model updates. Therefore, even if some clients fail to send model updates promptly due to various reasons, the global model on the server can still be updated by most clients, resulting in a usable global model. To further illustrate this point, we sequentially aggregated local model updates from client 1 to client 10 in one-shot federated fine-tuning of Llama-7b on the Wizard dataset. We tested the global model’s performance on the MT-bench as we aggregated updates from 1, 2, 3, and up to 10 clients. The results are displayed in Fig. 8. The model score increases as more clients contribute their local updates to the global model, indicating that each individual local model update provides an immediate improvement in global model performance. The red dashed line represents the model score in the synchronous FL setting, which matches the score of aggregating ten clients in the asynchronous setting. This behavior enables flexible participation under heterogeneous compute and partial availability, reduces server idle time, and better utilizes short or sporadic connectivity, which is often the case for edge devices and inter-organization deployments. From a systems perspective, it also simplifies recovery from stragglers and node restarts, since the aggregation state can progress whenever any subset of clients completes.
One-Shot Federated Fine-Tuning Naturally Mitigates Client-Side Privacy Threats
In a traditional FL algorithm, clients repeatedly receive new global model parameters each round, which could lead to client-side privacy issues. Malicious clients can exploit model inversion [fredrikson2015model, zhang2020secret] and gradient inversion attacks [huang2021evaluating] to recover private training samples or user inputs from other clients [wei2023client]. These attacks heavily rely on access to the global model parameters and certain data distribution information. However, in one-shot FL, the server can choose not to send back global parameters to the clients and only provide an API of the fine-tuned model. By doing this, it can eliminate the possibility of client-side privacy attacks. This design shifts the trust boundary from repeated parameter release to a service interface with authentication, access control, and logging, which is easier to monitor and audit. It is also compatible with defenses such as secure aggregation and client-side differential privacy, which are orthogonal to the one-shot protocol and can be enabled without changing the communication pattern. We note that while this reduces a class of client-side threats, it does not by itself address server-side risks such as model extraction; appropriate rate limiting, watermarking, and usage policies remain necessary in deployments that expose a public or partner-facing API.
VI Related Work
One-Shot Federated Learning
One-shot federated learning refers to learning the parameters of the global model in a single round of communication between clients and the server [guha2019one, liu2025one, guan2025capture, amato2025towards]. There are two main strategies for optimizing one-shot FL, neuron matching and knowledge distillation. Neuron matching is based on the permutation symmetry of neural networks [ainsworth2022git], which means that client model parameters can be aligned according to a common ordering and then be averaged. Previous works use algorithms such as the Fisher information matrix [jhunjhunwala2024fedfisher] and permutation matrix [wang2020federated] to match the local model parameters. The knowledge distillation methods aim at distilling knowledge from well-trained local models through public data [gong2021ensemble, li2020practical, heinbaugh2022data]. Some works also use distilled data to transfer knowledge between clients and the server [zhou2020distilled]. Recent works adopt generative models to help generate substitute data for the local dataset on the server [yang2024exploring, zhang2022dense]. These lines of work primarily seek to improve the protocol or optimization of one-shot FL (often for small or medium models), whereas we empirically discover and theoretically explain that one-shot FL already performs well for FMs. Hence we do not provide exhaustive head-to-head comparisons with every algorithmic variant; our contribution is to document the phenomenon and to offer an error analysis, rather than to introduce a new one-shot algorithm. Importantly, our findings are compatible with these techniques: neuron matching or distillation can be plugged into our setting, and our theory clarifies when and why such one-shot procedures should be effective in FMs.
Federated Fine-Tuning
Federated fine-tuning [cheng2021fine, orescanin2021federated] aims to adapt foundation models (FMs) on cross-domain, on-device datasets while preserving data privacy. To reduce communication and computation, many works employ PEFT methods such as LoRA [hu2021lora] within the federated setting [zhang2024towards]. As in classical FL, federated fine-tuning must contend with non-IID data [wang2024flora, cho2024heterogeneous] and personalization [wagner2024personalized]. In practice, the main deployment bottlenecks are the computational and bandwidth constraints of edge devices [woisetschlager2024federated, wang2025mobilea3gent, zheng2023autofed], for which one-shot FL can alleviate part of the cost.
While most federated fine-tuning studies on FMs pragmatically adopt fewer communication rounds and shorter local training for efficiency, prior work has not explicitly posed the difference between FMs and small models as a research question. Our work fills this gap: we empirically identify and theoretically explain why one-shot FL tends to perform well specifically for FMs, thereby offering a new lens on federated fine-tuning. Moreover, our findings are closely related to task arithmetic, task vectors, and model merging [hendel2023context, ilharco2022editing, zhao2024texttt], lines of work that have largely flourished in the FM era. This coincidence suggests a unifying hypothesis: the same factors that make one-shot FL effective for FMs (e.g., smoother landscapes, smaller update magnitudes, shorter effective horizons) may also underlie the empirical success of model merging and task arithmetic. These observations motivate a broader research agenda that explains and predicts the divergent behaviors of FMs versus small models in federated fine-tuning.
VII Conclusion
In this paper, we tackle the critical issue of high communication costs that limit the practical application of federated fine-tuning. Through a series of experiments, we demonstrate that multi-round communication is not necessary for fine-tuning FMs, as one-shot federated fine-tuning achieves comparable performance. We then provide a theoretical analysis to explain why one-shot federated fine-tuning is effective for large FMs and validate our findings with empirical evidence. Our extensive experiments show that one-shot federated fine-tuning performs on par with multi-round federated fine-tuning across 5 different models and 3 diverse tasks. This method significantly reduces communication overhead, making federated fine-tuning more feasible and efficient, especially for large-scale models. Moreover, one-shot federated fine-tuning supports asynchronous local updates and enhances security by minimizing data exposure during the training process.
Although this work does not propose a new algorithm, it offers a principled explanation for why one-shot FL can work well in large models. The analysis clarifies the conditions under which the one-shot gap remains small and provides insight into related phenomena such as task vectors and model editing, where small, targeted updates also transfer effectively in large models. Importantly, the empirical results align with the theoretical predictions and show that the bounds and assumptions connect with practical training behavior rather than remaining purely abstract. These findings make it possible to harness the power of large FMs in environments with limited communication resources, thereby broadening the accessibility and utility of advanced AI technologies. We view this as an initial step toward a more theory-grounded understanding of communication-efficient adaptation in federated settings, and we expect the framework to inform future studies on robust aggregation in FL under realistic system constraints.
[Uncaptioned image] Ziyao Wang is a third-year Ph.D. student in the Department of Electrical and Computer Engineering at the University of Maryland, College Park, advised by Prof. Ang Li. He received his B.E. degree from Wuhan University, China. His research focuses on improving the efficiency and trustworthiness of collaborative LLM systems, with interests in areas such as federated learning and server-device collaboration. He is also broadly interested in machine learning security and privacy. His research have been published in top conferences and journals such as NeurIPS, ICML, ICLR, EMNLP, CCS, TIFS, and TDSC. He has served as a reviewer for top-tier conferences (NeurIPS, ICML, ICLR, EMNLP) and journals (TMC, IEEE IoT).
[Uncaptioned image] Bowei Tian is a second-year Ph.D. student in the Department of Electrical and Computer Engineering at the University of Maryland, College Park, advised by Prof. Ang Li. He received his B.E. degree from Wuhan University, China. His research interests include causal reasoning and representation learning, AI security and privacy, and computer vision. His work has been published in top conferences and journals such as NeurIPS, ICLR, EMNLP, Ubicomp, ECCV, and TDSC. He has also served as a reviewer for top-tier conferences, including NeurIPS and AAAI.
[Uncaptioned image] Yexiao He is a third-year Ph.D. student in the Department of Electrical and Computer Engineering at the University of Maryland, College Park, advised by Prof. Ang Li. He received his B.E. and M.E. degrees from the University of Electronic Science and Technology of China. His research focuses on LLMs, AI for healthcare, and neuro-symbolic AI. His work has been published in top conferences such as NeurIPS and MobiSys.
[Uncaptioned image] Zheyu Shen is a third-year Ph.D. student in the Department of Electrical and Computer Engineering at the University of Maryland, College Park, advised by Prof. Ang Li. He received his M.S. degree in Computer Science from the University of Southern California and his B.E. degree from Northwestern Polytechnical University, China. His research focuses on machine learning systems, with particular interests in efficient LLM training and serving, federated learning, and distributed systems. He is also broadly interested in topics at the intersection of machine learning and edge computing. His work has been published at top conferences such as NeurIPS, EMNLP and MobiSys. He has also served as a reviewer for ICLR and NeurIPS.
[Uncaptioned image] Guoheng Sun is a second-year Ph.D. student in the Department of Electrical and Computer Engineering at the University of Maryland, College Park, advised by Prof. Ang Li. He received his B.E. degree from Sichuan University, China. His research interests lie in improving the efficiency and trustworthiness of LLMs, particularly through training-free or training-less approaches. He is also broadly interested in privacy, safety alignment, and the pretraining process of LLMs. He has serves(d) as a reviewer for top-tier conferences such as NeurIPS and ICLR. He is one of the awardees of the 2025 North America Qualcomm Innovation Fellowship.
[Uncaptioned image] Yuhan Liu is a first-year Ph.D. student in the School of Electronic Engineering and Computer Science at Queen Mary University of London, affiliated with the Centre for Digital Music. She received her Bachelor’s and Master’s degrees from Beijing Institute of Technology, China. Her research interests focus on machine learning for music analysis, specifically music source separation and music understanding.
[Uncaptioned image] Luyang Liu is a Research Scientist at Google DeepMind, focusing on foundation models, representation learning, and federated learning. He is a core contributor to Gemini models, with a focus on its thinking capability. He is also interested in enabling LLM with new capabilities, including latent thinking and text diffusions (i.e., Gemini Diffusion). Before that, he worked on representation learning, federated learning, and health research at Google Research. Before joining Google, Luyang conducted research in mobile/edge computing and road safety during his Ph.D. at Rutgers. His works have been published in top-tier venues such as ICML, NeurIPS, AAAI, MobiCom, MobiSys, Nature, and Science.
[Uncaptioned image] Meng Liu received her M.S. degree from University of Maryland, College Park, and B.S. from Henan University, China. Her research interests focus on interpretable machine learning, efficient machine learning, and AI for financial data.
[Uncaptioned image] Ang Li is an Assistant professor in the Department of Electrical and Computer Engineering at the University of Maryland College Park. Before joining UMD, he was a research associate at Qualcomm AI Research. He received Ph.D. in Electrical and Computer Engineering from Duke University in 2022. His research interests lie in the intersection of machine learning and edge computing, with a focus on building large-scale networked and efficient intelligent systems. His research has been recognized with a variety of awards, including the 2024 ACM CCS Distinguished Paper Award, 2022 IEEE TCCPS Outstanding Ph.D. Dissertation Award, 2022 Duke ECE Department Outstanding Dissertation Award, and ACM KDD Best Student Paper Award in 2020.